Pre-Lab Calculations 1. Use the Henderson-Hasselbalch equation to calculate the mass of solid sadium acetate required to mix with 50.0 mL of 0.10 M acetic acid to prepare a pH 4 buffer. Ka for acetic acid is 18 x 10 Use the Henderson-Hasselbalch equation to calculate the mass of solid ammonium chloride required to mix with 50.0 mL of 0.10 M ammonia to prepare a pH 10 buffer. The Kb for ammonia is 1.8 x 10 2. Purpose The purpose of this experiment is to explore the concept of equilibria in the context of acids and bases, to understand the nature of acid-base equilibrium, buffer solutions, and pH. Safety Precautions As with all experiments, you will be working in a potentially hazardous environment and must obey all lab safety rules. Hydrochloric acid, acetic acid, and sodium hydroxide are corrosive and will cause burns if spilled on skin or inhaled. Thi s experiment doe s not generate any hazardous waste. The solutions will be collected for pretreatment by your instructor. Materials To perform this experiment, you will need the following: Laboratory tape, la 0.1 M bels, or markers acid (HCI) Several small beakers beakers ah
Answer #1

POH: 14-PH = 14-10, 4 4 : 4.74 + log [4H23 [0-107 4-4.74- .. log Co 1o 0-74 Lo.10] 0.74 I0 N H 0.018197 4 0.018197 mol NH4 4Solu tion Henderson Hasselbach Eguation; base AcidJ Pkalog Kalog Cx 0) 474 8X10 [b a seJ 4 = 4.74 + log LbaseJ o.10] 4-4.74 0